OpenStudy (anonymous):

A diameter of 380 feet and a maximum depth of 50 feet. Find the equation of this parabola with equation y= ax^2.

OpenStudy (anonymous):

V=(0,50)

OpenStudy (anonymous):

by diameter do you mean latus rectum

OpenStudy (anonymous):

190=a x 50^2 i think since x=(-190, 50) and (190,50)

OpenStudy (anonymous):

kkkkkkkkkkk

OpenStudy (anonymous):

then i get y=19/250 = a?

OpenStudy (anonymous):

maybe x=190 and 50 is y so equation could be 50=a * 190^2

OpenStudy (anonymous):

http://www.purplemath.com/modules/parabola4.htm

OpenStudy (anonymous):

i think y=19/5 x^2 after reviewing purple math - does anyone agree?

OpenStudy (anonymous):

yes , i agree well a medal would be gr8 for me

OpenStudy (anonymous):

yes , i agree well a medal would be gr8 for me